A 1,200Kg boat is sunk in the harbor and sits on the bottom a depth of 30m. A plan to raise the boat using ping pong balls has been put in place. The boat is sealed so that the ping pong balls can be added to the boat and will remain in the boat displacing enough water to raise it to the surface. An average ping pong ball is 4 cm in diameter.

(a) Assume that all of the ping pong balls a perfectly sealed so that they will not take on water and that they are essentially massless. How many ping pong balls are needed to raise the boat?

(b) Now let’s make it more complicated Ping Pong balls actually have a typical mass of 2.7g and 0.2% of them are poorly made and will take on water. If add these considerations, (mass of the ping pong balls and % that leak), how many will you need?
Answer #1

a)   The boat is sealed so that the ping pong balls can be added to the boat and will remain in the boat displacing enough water to raise it to the surface.

weight of the boat =buoyancy force

1200*9.8 = (4/3)*pi*r^3 * n* 1*9.8 ( n= no. of ball)

2.68*10^-4*n =1200

n= 4471483

b)

1200*9.8+ n*0.998*2.7*10^-3*9.8 = (4/3)*pi*r^3 * n*0.998* 1*9.8

1200+ 2.6946*10^-3*n=2.67*10^-4*n

we will not get any valid soln.